a carpenter uses a 12 foot timber to brace a wall frame 8 feet high. if the construction site is level what is the distance the timber will extend out from the wall frame ?
Answer by algebrahouse.com(1659)

Pythagorean Theorem
a² + b² = c²
a and b are the legs
c is the hypotenuse {the side across from the right angle}

12 is the hypotenuse
8 is one of the legs
b is the distance from the wall

8² + b² = 12² {substituted into Pythagorean Theorem}
64 + b² = 144 {evaluated exponents}
b² = 80 {subtracted 64 from each side}
b = √80 {took square root of each side}
b = √16 √5 {broke down radical}
b = 4√5 {evaluated square root of 16
or
b ≈ 8.9 {decimal form}

To ask Algebra House, visit: www.algebrahouse.com